OpenStudy (anonymous):

A 1600 kg car on a flat ground is moving 6.25 m/s. Its engine creates a 1150N forward force as the car moves 45.8m. What is its final KE?

OpenStudy (anonymous):

\[intial KE=\frac{ 1 }{2 }mv ^{2} = \frac{ 1 }{2 }*1600*6.25^{2}=31250J\]\[Enginework=F*d=1150*45.8=52670J\]\[finalKE=initialKE+Enginework=31250+52670=83920J\]

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

How about: A 1210 kg rollercoaster car is moving 8.31m/s. Before entering the station, it rolls up a 2.26m hill. How fast is it moving then?

OpenStudy (anonymous):

this time first calculate car's initial KE and substract the PE required for 2.26m rise from it. this will give you present KE. from that calculate present speed... :)

OpenStudy (anonymous):

how do you calculate the present speed?